OpenStudy (anonymous):

A manufacturer finds that the revenue generated by selling x units of a certain commodity is given by the function R(x) = 60x − 0.2x2, where the revenue R(x) is measured in dollars. What is the maximum revenue, and how many units should be manufactured to obtain this maximum?

OpenStudy (anonymous):

i don't know, but would you differentiate once to make it 60 ?

OpenStudy (goformit100):

OK

OpenStudy (anonymous):

Im erally not sure what to do

OpenStudy (anonymous):

ya, won't you just differentiate?

OpenStudy (anonymous):

I am not sure, but I guess this might be correct, R(x) = 60x - 0.2x^2 Differentiating w.r.t x; R` = 60 - 0.4x equate to zero, 60 - 0.4x = 0 x= 150 therefore max. revenue will be, 8940. (substituted value of x in first eq.)

OpenStudy (anonymous):

the rev was wrong but the units were right

OpenStudy (anonymous):

ya, revenue will be 4500 . my bad

OpenStudy (anonymous):

lol yeah that was it. thank you so much